# with Onset-Aware Conditioning for Synchronized Video-to-Audio Synthesis

### v2

@@ 1 Introduction
Existing video-to-audio approaches naively learn latent audio representation through reconstruction loss, which can lead to suboptimal quality. Recently, REPA [39] has attempted to leverage pretrained features [24] as a constraint to guide latent representations in image synthesis. Inspired by this, we explore applying a similar constraint-driven approach to latent audio representation learning. Directly applying methods like REPA is non-trivial, and we carefully design TRA by (1) dynamically adjusting the alignment strength based on the diffusion noise schedule, (2) addressing sequence length mismatch, and (3) selecting appropriate audio encoder outputs. This prevents over-regularization while preserving semantic grounding throughout training. Yet adapting this approach for video-to-audio synthesis is challenging, as imposing strict constraints on latent representations in early timesteps, when noise dominates, can hinder learning rather than improve alignment. Instead, a more adaptive strategy is required to gradually refine alignment as the latent representation evolves. Additionally, achieving precise synchronization at the event level remains an open challenge. Existing methods struggle to capture fine-grained auditory cues tied to specific visual events, often generating misaligned or unnatural sounds. This issue arises from the lack of explicit mechanisms to detect and condition on key auditory events such as footsteps, object interactions, or percussive impacts, which serve as strong temporal anchors for realistic audio synthesis.
@@ 2 Related Works
Video-to-audio generation. Video-to-audio generation aims to synthesize temporally synchronized and semantically relevant audio from video. Transformer-based autoregressive methods like SpecVQGAN [13] offer flexibility but suffer from slow inference and weak alignment. Im2Wav [27] extends this to images via CLIP [26] embeddings but struggles with dynamic video content. Diffusion-based models [12, 28, 37, 36, 38] improve quality but require multiple steps and large parameters. Diff-Foley [22] enhances alignment via latent diffusion, while See & Hear [34] improves multimodal integration but lacks fine-grained temporal precision. FoleyCrafter [40] incorporates a temporal controller for high-quality Foley sound but remains computationally heavy. Frieren [31] reduces steps via rectified flow matching but may miss sparse, event-driven dynamics. MDSGen [25] employs masked diffusion for efficiency but still relies on multiple steps for fidelity. Furthermore, several works incorporate auxiliary temporal features as condition information to guide the generation process. TiVA [30] leverages smoothed energy control, and ReWaS [14] uses down-sampled mel-spectrograms. However, these signals often introduce temporal redundancy or smoothing artifacts, making them less effective at capturing fine-grained event boundaries critical for precise alignment. Our TARO framework addresses these issues using timestep-adaptive alignment, and onset-aware conditioning to enhance audio quality, synchronization, and computational efficiency.
@@
Audio Representation Learning. Recent advances in audio representation learning have significantly improved the ability to extract meaningful features for audio understanding, offering valuable insights for generative models. Self-supervised learning frameworks such as wav2vec 2.0 [2] capture robust speech representations, while CLAP [33] employs contrastive language-audio pretraining to learn general-purpose audio semantics. Transformer-based models like BEATs [5] leverage masked spectrogram learning for audio classification, whereas EAT [6] introduces an efficient audio transformer trained with utterance-frame objectives. Despite their effectiveness, these models have yet to be fully leveraged for video-to-audio tasks, where precise temporal alignment and semantic consistency are critical. Our approach integrates pretrained audio encoders into a timestep-adaptive alignment mechanism, ensuring that latent representations evolve meaningfully across synthesis.
@@ 4.1 Overview of TARO
To enable effective multimodal integration, our framework concatenates the projected visual features and audio latent embeddings along the sequence dimension before feeding them into the transformer blocks, as illustrated in Figure 4. Given the differences in feature spaces and numerical scales between these modalities, Adaptive Layer Normalization (AdaLN) is applied separately to each modality, ensuring consistent processing and preventing imbalance between visual and audio representations. These processed tokens then interact through a joint attention mechanism, where query, key, and value representations from both modalities are concatenated and jointly attended, allowing the model to capture fine-grained audio-visual correspondences and refine the latent representation synchronized. A key component of our framework is the Timestep-Adaptive Representation Alignment (TRA), which dynamically aligns latent representations with high-quality pretrained audio encoders [6, 5, 33, 2]. This adaptive strategy leverages pretrained audio priors to enhance both audio fidelity and synchronization with visual content.
@@ 4.3 Time-Adaptive Representation Alignment
The standard alignment method [39] apply fixed constraints throughout synthesis, which can hinder learning in early timesteps when noise dominates. TRA dynamically adjusts alignment strength based on the noise schedule, allowing latent representations to evolve progressively. This adaptive approach prevents overly strict constraints in noisy stages while gradually reinforcing structure as synthesis progresses, ensuring smooth and effective alignment. Given a pretrained encoder $f_{enc}$ and a clean audio sample $x_{a}$, we compute its embedding as $y_{a}=f_{enc}(x_{a})\in\mathbb{R}^{L_{a}\times D_{a}}$, where $L_{a}$ represents the number of patches and $D_{a}$ denotes the embedding dimension. To align with this reference, we use a trainable projection head $h_{\phi}$, which transforms the model latent representation $z_{t}\in\mathbb{R}^{L_{z}\times D_{z}}$ at timestep $t$ into the same $y_{a}$ feature space as $h_{\phi}(z_{t})\in\mathbb{R}^{L_{z}\times D_{a}}$. This projection ensures that the learned representation matches the structured features of clean audio, enabling effective alignment. To achieve this alignment, we introduce a time-weighted alignment loss, adjusting alignment strength dynamically based on the noise schedule:
@@
Figure 4: One Multimodal Transformer Blocks. The block integrates onset cues, visual features, and latent audio representations through adaptive modulation and joint attention.$+$ denotes summation, and C represents concatenation.
@@
[…]
@@ 5.2 Main Results
Fig. 5 presents the qualitative comparison of our TARO and SOTA video-to-audio methods [13, 27, 22, 34, 40, 31, 25]. TARO exhibits superior audio quality and synchronization, closely aligning generated spectrograms with ground-truth. On the right, it effectively reproduces two distinct crowing events in a pheasant video, ensuring seamless alignment with visual cues and enhancing perceptual coherence. These results highlight TARO’s ability to model fine-grained audio-visual correlations more effectively than prior methods. Additional audio samples and visualizations are provided in Suppl Sec. F.
@@ 5.3 Ablation Studies
As shown in Table 3, removing both components leads to the lowest scores. TRA alone improves distribution matching (FD, FAD), while OAC enhances synchronization (Acc, CLIP Score). Combining both achieves the best results, demonstrating their complementary effects on audio quality and alignment. Fig. 6 visualizes the impact of OAC. Without it, the generated spectrogram lacks a clear temporal structure and deviates from the ground truth. In contrast, incorporating OAC, our model closely aligns with the ground truth, ensuring precise synchronization between generated audio and visual events. Full metric evaluations for all ablation are provided in Suppl Sec. C.
@@ Appendix D Additional Results & Analysis
Training efficiency As shown in Fig. 7, TRA improves both training stability and generation quality compared to the vanilla model.

### v3

@@ 1 Introduction
Existing video-to-audio approaches naively learn latent audio representation through reconstruction loss, which can lead to suboptimal quality. Recently, REPA [44] has attempted to leverage pretrained features [28] as a constraint to guide latent representations in image synthesis. Inspired by this, we explore applying a similar constraint-driven approach to latent audio representation learning. Directly applying methods like REPA is non-trivial, so we carefully design TRA by (1) dynamically adjusting the alignment strength based on the diffusion noise schedule, (2) addressing sequence length mismatch, and (3) selecting appropriate audio encoder outputs. This prevents over-regularization while preserving semantic grounding throughout training. In particular, adapting this regularization for video-to-audio synthesis is challenging, as imposing strict constraints on latent representations in early timesteps, when noise dominates, can hinder learning rather than improve alignment. Instead, a more adaptive strategy is required to gradually refine alignment as the latent representation evolves. Additionally, achieving precise synchronization at the event level remains an open challenge. Existing methods struggle to capture fine-grained auditory cues tied to specific visual events, often generating misaligned or unnatural sounds. This issue arises from the lack of explicit mechanisms to detect and condition on key auditory events such as footsteps, object interactions, or percussive impacts, which serve as strong temporal anchors for realistic audio synthesis.
@@ 2 Related Works
Video-to-audio generation. Video-to-audio generation, a core task in multimodal generation [31, 26], aims to synthesize temporally synchronized and semantically relevant audio from video. Transformer-based autoregressive methods like SpecVQGAN [15] offer flexibility but suffer from slow inference and weak alignment. Im2Wav [32] extends this to images via CLIP [30] embeddings but struggles with dynamic video content. Diffusion-based models [13, 33, 42, 41, 43, 14] improve quality but require multiple steps and large parameters. Diff-Foley [24] enhances alignment via latent diffusion, while See & Hear [39] improves multimodal integration but lacks fine-grained temporal precision. FoleyCrafter [45] incorporates a temporal controller for high-quality Foley sound but remains computationally heavy. Frieren [36] reduces steps via rectified flow matching but may miss sparse, event-driven dynamics. MDSGen [29] employs masked diffusion for efficiency but still relies on multiple steps for fidelity. Furthermore, several works incorporate auxiliary temporal features as condition information to guide the generation process. TiVA [35] leverages smoothed energy control, and ReWaS [16] uses down-sampled mel-spectrograms. However, these signals often introduce temporal redundancy or smoothing artifacts, making them less effective at capturing fine-grained event boundaries critical for precise alignment. Our TARO framework addresses these issues using timestep-adaptive alignment, and onset-aware conditioning to enhance audio quality, synchronization, and computational efficiency.
@@
Audio Representation Learning. Recent advances in representation learning [11, 25] have significantly improved the ability to extract meaningful features for audio understanding, offering valuable insights for generative models. Self-supervised learning frameworks such as wav2vec 2.0 [2] capture robust speech representations, while CLAP [38] employs contrastive language-audio pretraining to learn general-purpose audio semantics. Transformer-based models like BEATs [5] leverage masked spectrogram learning for audio classification, whereas EAT [6] introduces an efficient audio transformer trained with utterance-frame objectives. Despite their effectiveness, these models have yet to be fully leveraged for video-to-audio tasks, where precise temporal alignment and semantic consistency are critical. Our approach integrates pretrained audio encoders into a timestep-adaptive alignment mechanism, ensuring that latent representations evolve meaningfully across synthesis.
@@ 4.1 Overview of TARO
To enable effective multimodal integration, our framework concatenates the projected visual features and audio latent embeddings along the sequence dimension before feeding them into the transformer blocks, as illustrated in Figure 4. Given the differences in feature spaces and numerical scales between these modalities, Adaptive Layer Normalization (AdaLN) is applied separately to each modality, ensuring consistent processing and preventing imbalance between visual and audio representations. These processed tokens then interact through a joint attention mechanism, where query, key, and value representations from both modalities are concatenated and jointly attended, allowing the model to capture fine-grained audio-visual correspondences and refine the latent representation. A key component of our framework is the Timestep-Adaptive Representation Alignment (TRA), which dynamically aligns latent representations with high-quality pretrained audio encoders [6, 5, 38, 2]. This adaptive strategy leverages pretrained audio priors to enhance both audio fidelity and synchronization with visual content.
@@ 4.3 Time-Adaptive Representation Alignment
The standard alignment method [44] applies fixed constraints throughout synthesis, which can hinder learning in early timesteps when noise dominates. TRA dynamically adjusts alignment strength based on the noise schedule, allowing latent representations to evolve progressively. This adaptive approach prevents overly strict constraints in noisy stages while gradually reinforcing structure as synthesis progresses, ensuring smooth and effective alignment. Given a pretrained encoder $f_{enc}$ and a clean audio sample $x_{a}$, we compute its embedding as $y_{a}=f_{enc}(x_{a})\in\mathbb{R}^{L_{a}\times D_{a}}$, where $L_{a}$ represents the number of patches and $D_{a}$ denotes the embedding dimension. To align with this reference, we use a trainable projection head $h_{\phi}$, which transforms the model latent representation $z_{t}\in\mathbb{R}^{L_{z}\times D_{z}}$ at timestep $t$ into the same $y_{a}$ feature space as $h_{\phi}(z_{t})\in\mathbb{R}^{L_{z}\times D_{a}}$. This projection ensures that the learned representation matches the structured features of clean audio, enabling effective alignment. To achieve this alignment, we introduce a time-weighted alignment loss, adjusting alignment strength dynamically based on the noise schedule:
@@
Figure 4: One Multimodal Transformer Blocks. The block integrates onset cues, visual features, and latent audio representations through adaptive modulation and joint attention.++ denotes summation, and C represents concatenation.
@@
Table 1: Quantitative comparisons of video-to-audio models on the VGGSound dataset. We use colors to denote the first, second and third places respectively. $\dagger$}: MOS-AQ and MOS-AV scores are evaluated on samples from both VGGSound and Landscape datasets.
@@ 5.2 Main Results
Figure 5 presents the qualitative comparison of our TARO and SOTA video-to-audio methods [15, 32, 24, 39, 45, 36, 29]. TARO exhibits superior audio quality and synchronization, closely aligning generated spectrograms with ground-truth. On the right, it effectively reproduces two distinct crowing events in a pheasant video, ensuring seamless alignment with visual cues and enhancing perceptual coherence. These results highlight TARO’s ability to model fine-grained audio-visual correlations more effectively than prior methods. Additional audio samples and visualizations are provided in Suppl Sec. F.
@@ 5.3 Ablation Studies
As shown in Table 3, removing both components leads to the lowest scores. TRA alone improves distribution matching (FD, FAD), while OAC enhances synchronization (Acc, CLIP Score). Combining both achieves the best results, demonstrating their complementary effects on audio quality and alignment. Figure 6 visualizes the impact of OAC. Without it, the generated spectrogram lacks a clear temporal structure and deviates from the ground truth. In contrast, incorporating OAC, our model closely aligns with the ground truth, ensuring precise synchronization between generated audio and visual events. Full metric evaluations for all ablation are provided in Suppl Sec. C.
@@ 6 Conclusion
Acknowledgments
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [RS-2021-II212068, Artificial Intelligence Innovation Hub (Seoul National University)] and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea) & Gwangju Metropolitan City.
@@ Appendix D Additional Results & Analysis
Training efficiency As shown in Figure 7, TRA improves both training stability and generation quality compared to the vanilla model.